# Whitney Approximation: domains and bounds

### v1

@@ 5. Whitney Approximation with Bounds
(5.1) $42\delta_{n+1}\leqslant\delta_{n},\qquad\sum_{m=n}^{\infty}\delta_{m}M_{m+1}\leqslant\varepsilon_{n}/$ 4 .
@@
Since $f\neq 0$ by assumption, we can take $n_{0}$ with $G_{0n_{0}}>0$. Then with $c:=\varepsilon_{0}^{3}/G_{0n_{0}}^{2}$ we have $c\lambda_{n}\geqslant n^{5}G_{0n}$ for each $n$. The function $t\mapsto t^{2}\operatorname{e}^{-t}\colon{\mathbb{R}}^{>}\to{\mathbb{R}}$ takes on its maximum value $14\operatorname{e}^{-2}<$ 1 at $t=2$, so $t\operatorname{e}^{-t}<t^{-1}$ for each $t>0$. Hence for $n\geqslant 1$:
@@
Let $f$, $\varepsilon$, $\rho$ be as in the statement of Theorem 2. We choose $a_{n}$$b_{n}$ as in (${\mathbb{R}}$) above, with $\delta:=\sqrt{2}$, and set $\varepsilon_{n}:=\varepsilon(b_{n+1})={\varepsilon\big{(}\sqrt{2}(n+1)\big{)}}$ and $r_{n}:=\lfloor\rho(b_{n+1})\rfloor$. Then conditions (i)–(v) hold, with (iv) a consequence of the convexity of $\varepsilon$. We also have $\rho_{n}=1$ and $U={\mathbb{C}}$ by (5.4 ${\mathbb{R}}$). We now construct the real entire function $\widehat{g}$ and its restriction $g\in\mathcal{C}^{\omega}({\mathbb{R}})$ as described above, in the process choosing the quantities $D_{n}$, $\delta_{n}$, $\lambda_{n}$, $c_{n}$ as indicated, in the case (${\mathbb{R}}$). (In particular, $c_{n}=c_{*}/n^{2}$ for $n\geqslant 1$, where the constant $c_{*}>0$ doesn’t depend on $\rho$.) Then for $t\in L_{n}$ and $k\leqslant\rho(\lvert t\rvert)$ we have $\lvert t\rvert\leqslant b_{n+1}$, thus $k\leqslant r_{n}$ and so
$\lvert(f-g)^{(k)}(t)\rvert\leqslant\lVert f-g\rVert_{L_{n};\,r_{n}}<\varepsilon_{n}=\varepsilon(b_{n+1})\leqslant\varepsilon(\lvert t\rvert).$
Thus $\lvert(f-g)^{(k)}(t)\rvert<\varepsilon(\lvert t\rvert)$ for all $t$ and $k\leqslant\rho(\lvert t\rvert)$. We now aim to estimate $\lvert\widehat{g}(z)\rvert$ when $\lvert z\rvert\leqslant t$. We need two lemmas. In the first one we bound
@@
As we are in case (${\mathbb{R}}$), we can assume that we have $c,d\in{\mathbb{R}}^{\geqslant 1}$ (independent of $f$, $\varepsilon$, $\rho$) such that $D_{mn}=cm^{dm}$ for all $m$$n$. We have
@@
[…]
@@
We now construct the holomorphic function $\widehat{g}\colon U\to{\mathbb{C}}$ with $\widehat{g}({\mathbb{R}}^{>})\subseteq{\mathbb{R}}$ as described before, and set $g:=\widehat{g}|_{{\mathbb{R}}^{>}}\in\mathcal{C}^{\omega}({\mathbb{R}}^{>})$. As in the proof of Theorem 2, we see that for $t\in L_{n}$ and $k\leqslant\rho(t)$ we have $\lvert(f-g)^{(k)}\rvert<\varepsilon(t)$. Since $\bigcup_{n}L_{n}={\mathbb{R}}^{>}$, this yields $\lvert(f-g)^{(k)}(t)\rvert<\varepsilon(t)$ for all $t>0$ and $k\leqslant\rho(t)$. Now put $V_{n}:=U_{n}\cap V$, so $V=\bigcup_{n}V_{n}$, and set $V_{-1}:=\emptyset$.
@@
[…]

### v2

@@ 5. Whitney Approximation with Bounds
(5.1) $2\delta_{n+1}\leqslant\delta_{n},\qquad\sum_{m=n}^{\infty}\delta_{m}M_{m+1}\leqslant\varepsilon_{n}/4.$
@@
Since $f\neq 0$ by assumption, we can take $n_{0}$ with $G_{0n_{0}}>0$. Then with $c:=\varepsilon_{0}^{3}/G_{0n_{0}}^{2}$ we have $c\lambda_{n}\geqslant n^{5}G_{0n}$ for each $n$. The function $t\mapsto t^{2}\operatorname{e}^{-t}\colon{\mathbb{R}}^{>}\to{\mathbb{R}}$ takes on its maximum value $4\operatorname{e}^{-2}<1$ at $t=2$, so $t\operatorname{e}^{-t}<t^{-1}$ for each $t>0$. Hence for $n\geqslant 1$:
@@
[…]
@@
As we are in case ($\mathbb{R}$}), we can assume that we have $c,d\in{\mathbb{R}}^{\geqslant 1}$ (independent of $f$, $\varepsilon$, $\rho$) such that $D_{mn}=cm^{dm}$ for all mmnn. We have
@@
$\delta_{n}^{-1}=\frac{c2^{n+3}(r_{n+1})^{dr_{n+1}}}{\varepsilon_{n}-\varepsilon_{n+1}}\leqslant\frac{2c\,2^{s/\sqrt{2}}\rho(s)^{d\rho(s)}}{\Delta\varepsilon(s)}$
@@
[…]
@@
$\delta_{n}^{-1}\leqslant\frac{c2^{s/\delta}\big{(}\rho(s)(s/\delta-2)\big{)}^{d\rho(s)}}{\varepsilon_{n}-\varepsilon_{n+1}}\leqslant\frac{(c/\delta)\,2^{s/\delta}\,\big{(}\rho(s)(s/\delta)\big{)}^{d\rho(s)}}{\Delta\varepsilon(s)},$